# Relationship Discovery for Drug Recommendation

### v1

@@ 5.4 Parameters Analysis
5.5 Quality Study
Figure 4: The performance of co-occurrence-based and causality-based methods on Jaccard.
To validate the significance of introducing causality, we compare the performance of causal relationships and co-occurrence relationships in our model on MIMIC-III. We create a comparative model by replacing all causal relationships in DisMed with co-occurrence relationships. Specifically, we remove causal inference methods and replace the causal effects between diseases and medications in treatment analysis with co-occurrence rates. Since co-occurrence methods cannot capture the pathological relationships between conditions, we remove an entire process of condition analysis and fuse them using equal-weight addition. The subsequent recommendation methods remain the same as the original model.
@@ 5.5 Quality Study
5.6 Case Study
(a) Quantitative relationship between specific dis/proc and med.
(b) Prescription principles for a specific visit.
Figure 5: The learning process under co-occurrence-based and causality-based approaches from the real clinical sample, where $d_{1}$ is Gastrointest hemorr, $d_{2}$ is Candidal esophagitis, $d_{3}$ is B complex deficiency, $d_{4}$ is Regional enteritis, $p_{1}$ is Sm bowel endoscopy, $m_{1}$ is Drug for peptic ulcer and GORD, $m_{2}$ is Antifungals for topical use, $m_{3}$ is Other mineral supplements, $m_{4}$ is Hypnotics and sedatives, $m_{5}$ is Other analgesics and antipyretics, $m_{6}$ is Intestinal antiinflammatory agents and $m_{7}$ is Sulfonamides and trimethoprim.
@@ 5.6 Case Study
To more intuitively demonstrate how causality-based methods capture personalized patient representations, we employed the same co-occurrence-based comparative model as in the previous subsection and visualized the learning process of a real case under both approaches. Figure 5 is derived from a clinical visit associated with gastric bleeding. Sub-figure 5(a) illustrates the quantified relationships between diseases/procedures and medications, while Sub-figure 5(b) provides a visualization of the learning results.
The left half of the Sub-figure 5(a) illustrates the co-occurrence-based relationships (co-occurrence rates) between diseases/procedures and medications. Those relationships learned from such set-to-set association are ambiguous, with a single medication exhibiting similar correlations with all types of diseases/procedures. In particular, $m_{5}$ (other analgesics and antipyretics), $m_{6}$ (intestinal anti-inflammatory agents), and $m_{7}$ (sulfonamides and trimethoprim) are frequently prescribed medications, each exhibiting high co-occurrence rates with a wide array of diseases. Meanwhile, $m_{3}$ (other mineral supplements), being relatively rare, has low co-occurrence rates with all diseases and procedures, making it challenging for the model to determine its specific association with any particular disease or procedure. Such misleading correlations, as depicted in the left half of the Sub-figure 5(b), result in co-occurrence-based prescription principles where popular medications appear to be associated with nearly every disease/procedure, whereas less common medications are unable to establish their appropriate relationships. This ultimately leads the model to preferentially recommend common medications in the prescription process, while overlooking rare medications. However, although this strategy is not rational, in certain cases, blindly recommending common medications can also yield correct outcomes. This can lead to positive feedback on such a strategy during training, resulting in the continual amplification of this error.
@@
In contrast, as shown in the right half of the Sub-figure 5(b), relationships based on causality (causal effects) are clearer. The point-to-point causal estimation allows each medication to be specifically correlated with its corresponding diseases or procedures. Therefore, we can address the spurious correlations arising from the common or rare occurrences of medications in co-occurrence-based relationships. Furthermore, as shown in the right half of the Sub-figure 5(b), by exploring the differences between patients’ conditions, we establish dynamic patient profiles. By integrating the precise relationships between conditions and medications, we ultimately construct more personalized patient representations. Our method learns that doctors only need to prescribe for $d_{4}$ (Regional enteritis) and $p_{1}$ (Small bowel endoscopy) related to $d_{1}$ (Gastrointest hemorr), without the need to prescribe specifically for $d_{1}$. In contrast, in the co-occurrence-based method, all these medications are prescribed for $d_{1}$.
@@
We demonstrate how causality-based methods enhance medication recommendation personalization and validate our proposed framework’s rationale via comparative analysis.
6 Conclusion
@@
We introduce a causality-based, condition-centric framework named DisMed, which effectively addresses the issues of ambiguous relationships between conditions and medications and the challenge of capturing differences among conditions, thereby significantly enhancing the personalization of patient representations. Extensive experiments on real-world datasets validate the efficacy of the proposed framework.

### v2

@@ 5.4. Parameters Analysis
5.5. Quality Study (RQ3)
(a) Results of the co-occurrence relationship-based model on the training and validation sets.
(b) Results of causality-based models on training and validation sets.
Figure 5. The performance of co-occurrence-based and causality-based methods on Jaccard.
[…]
To validate the significance of introducing causality, we compare the performance of causal relationships and co-occurrence relationships in our model. We create a comparative model by replacing all causal relationships in CausalMed with co-occurrence relationships. Specifically, we remove causal inference methods and replace the causal effects between diseases/procedures and medications with co-occurrence rates. Since co-occurrence methods cannot capture the pathological relationships between medical entities, we remove an entire process of health state learning and fuse them using equal-weight addition. The subsequent recommendation methods remain the same as the original model.
@@ 5.5. Quality Study (RQ3)
5.6. Case Study (RQ3&RQ4&RQ5)
(a) Quantitative relationships between specific diseases/adverse reactions and medications based on co-occurrence relationships.
(b) Quantitative relationships between specific diseases/adverse reactions and medications based on causal relationships.
(c) Prescription principles for a specific visit.
Figure 6. The learning process under co-occurrence-based and causality-based approaches from the real clinical sample.
[…]
@@ 5.6. Case Study (RQ3&RQ4&RQ5)
To more intuitively demonstrate how causality-based methods capture personalized patient representations, we employ the same co-occurrence-based comparative model as in the previous subsection and make transparent the learning process of patient representations in a real case under both approaches. This ultimately proves the rationality and interpretability of the recommended results in this study. Figure 6 is derived from a clinical visit associated with gastric bleeding. Sub-figure 6(a) and 6(b) illustrate the quantified relationships between diseases/procedures and medications, while Sub-figure 6(c) provides a visualization of the learning results.
Sub-figure 6(a) shows co-occurrence-based relationships between diseases/procedures and medications, which are ambiguous. Medications like $m_{5}$ (analgesics), $m_{6}$ (intestinal anti-inflammatory agents), and $m_{7}$ (sulfonamides) are frequently prescribed and show high co-occurrence rates with many diseases. In contrast, rare medications like $m_{3}$ (mineral supplements) have low co-occurrence rates, making specific associations difficult to determine. As depicted in the upper half of Sub-figure 6(c), this results in common medications being associated with nearly every disease/procedure while rare medications fail to establish appropriate relationships. Consequently, the model preferentially recommends common medications, overlooking rare ones, sometimes leading to positive feedback during training and amplifying the error.
@@
In contrast, relationships based on causality are clearer, as shown in Sub-figure 6(b). Point-to-point causal estimation links each medication to its specific diseases or procedures, avoiding spurious correlations found in co-occurrence-based relationships. The lower half of Sub-figure 6(c) shows that by analyzing differences in patients’ diseases/procedures, we create dynamic patient profiles. By integrating precise disease/procedure-medication relationships, we build more personalized patient representations. Our method learns that doctors only need to prescribe for $d_{4}$ (Regional enteritis) and $p_{1}$ (Small bowel endoscopy) related to $d_{1}$ (Gastrointest hemorr), without prescribing specifically for $d_{1}$. In contrast, the co-occurrence-based method prescribes all these medications for $d_{1}$.
@@
We demonstrate how causality-based methods enhance medication recommendation personalization and validate our proposed framework’s reasonability, transparency and interpretability via comparative analysis.
6. Conclusion
@@
We introduce a causality-based, patient health state-centric framework named CausalMed, which effectively addresses the issues of ambiguous relationships between diseases/procedures and medications and the challenge of capturing differences among diseases/ procedures, thereby significantly enhancing the personalization of patient representations. Extensive experiments on real-world datasets validate the efficacy of the proposed framework.
Acknowledgements.
This work was supported by the Innovation Capability Improvement Plan Project of Hebei Province (No. 22567637H), the S&T Program of Hebei(No. 236Z0302G), and HeBei Natural Science Foundation under Grant (No.G2021203010 & No.F2021203038).